# Axion dark matter from inflation-driven quantum phase transition

### 13 Feb 2024 v1

@@ Abstract
We propose a new mechanism to produce axion dark matter from inflationary fluctuations. Quantum fluctuations during inflation are strengthened by a coupling of the axion kinetic term to the inflaton, which we parametrize as an effective curvature $\kappa$ in the axion equation of motion. A nonvanishing curvature breaks the scale invariance of the axion power spectrum, driving a quantum phase transition with $\kappa$ as the order parameter. The axion power spectrum is proportional to the inverse comoving horizon to the power of $\kappa$. For positive $\kappa$ the spectrum gets a red tilt, leading to an exponential enhancement of the axion abundance as the comoving horizon shrinks during inflation. This enhancement allows sufficient axion production to comprise the entire dark matter relic abundance despite the ultralight mass. Our mechanism predicts a significantly different parameter space from the usual misalignment mechanism. It allows for axion-like particle dark matter with a much lower decay constant and thus a larger coupling to Standard Model particles. Much of the parameter space can be probed by future experiments including haloscopes, nuclear clocks, and CMB-S4. We can also generate heavier QCD axion dark matter than the misalignment mechanism.
@@ I Introduction
The usual misalignment mechanism for producing the axion relic abundance predicts a relationship between the axion mass and the PQ breaking scale Dine and Fischler (1983); Preskill et al. (1983); Arias et al. (2012). Assuming that the axion comprises all of the DM, this restricts the viable parameter space. For instance, for QCD axion DM, the mass should be of order $10^{-6}~{}{\rm eV}$ (assuming the initial misalignment angle is not fine-tuned) Grilli di Cortona et al. (2016). On the other hand, if we consider ALP DM lighter than $10^{-12}~{}{\rm eV}$, which is the regime where future quantum sensor probes are most sensitive Kim and Perez (2024); Flambaum and Samsonov (2023); Flambaum and Mansour (2023); Zhao et al. (2024), then the PQ breaking scale has to be higher than $10^{14}~{}{\rm GeV}$. Consequently, the couplings to Standard Model (SM) particles are highly suppressed and difficult to probe experimentally. This has motivated the development of new mechanisms for producing axion DM that allow for a different relationship between mass and coupling, such as nonstandard cosmologies Blinov et al. (2019), kinetic misalignment Co et al. (2020), ALP cogenesis Co et al. (2021), and trapped misalignment Di Luzio et al. (2021).
@@
In this work, we propose a new mechanism to produce axion DM from inflationary fluctuations, which allows for a much lower breaking scale than misalignment. The basic idea is to make use of quantum fluctuations during the cosmic inflationary epoch to efficiently produce the axion abundance. Typically, inflationary fluctuations do not lead to a sufficient DM production, due to the smallness of the axion mass Redi and Tesi (2023). However, if the axion kinetic term is coupled to the inflaton, the relic abundance can be exponentially enhanced [^1] A similar idea was first raised in Ref. Nakai et al. (2020) to produce dark photon DM through exponential enhancement, where a specific form of kinetic coupling was used.. The kinetic coupling breaks the scale-invariant power spectrum of a massless axion and, together with the quantum fluctuations during inflation, drives a “quantum phase transition” from the conformal symmetric phase to the broken phase. We will show how one can calculate the tilt in the power spectrum from the form of the kinetic coupling. We find that if the power spectrum is broken to a red tilt, the axion abundance gets exponentially enhanced by the number of inflationary e-folds, compensating the suppression of the DM density by the small axion mass.
@@
Our mechanism is applicable to both the QCD axion and ALPs. The relic density does not depend on the PQ breaking scale, which opens up a large window in the axion DM parameter space across a wide range of masses down to $\sim 10^{-24}~{}{\rm eV}$. The axion couplings can be much larger than one obtains in the misalignment mechanism. Much of the parameter space can be probed by future axion DM experiments such as haloscopes and nuclear clocks (see Fig. 2). Our mechanism can also potentially be probed by next-generation cosmic microwave background (CMB) experiments and $21$ cm telescopes.
@@ II.1 Framework
where $M_{\rm Pl}$ is the reduced Planck mass, $R$ is the Ricci scalar, and $V(\phi)$ is the inflaton potential. The axion $\eta$ is coupled to the inflaton $\phi$ through its kinetic term during inflation, with $K(\phi)$ an arbitrary function of the inflaton. The kinetic function $K(\phi)$ should smoothly reduce to unity at the end of inflation, such that the axion is decoupled from the inflaton thereafter and we are left with a canonical kinetic term. We take the usual flat FLRW metric $g_{\rm\mu\nu}=a^{2}(\tau){\rm diag}\left(-1,1,1,1\right)$, where $a$ is the scale factor and $\tau$ is the conformal time.
@@ II.2 Breaking conformal symmetry
where $H_{\rm inf}/2\pi$ is the Gibbons–Hawking temperature Gibbons and Hawking (1977) that characterizes the typical magnitude of quantum fluctuations during inflation. Note in the last step of Eq. (7) we have assumed $\left|\kappa\right|<1$.
@@
In particular, a positive $\kappa$ leads to a red tilt, with the axion power spectrum dominated by superhorizon modes. Note that in Eq. (7), we have $1/x=aH_{\rm inf}/k$, which means the power spectrum obeys a power law proportional to the inverse comoving horizon $aH_{\rm inf}$ to the power of $\kappa$. Since the comoving horizon is exponentially shrinking during inflation, the power spectrum will get exponentially enhanced if $\kappa>0$. For each mode $k$, it starts to grow only after exiting the horizon, i.e., $\left(aH_{\rm inf}\right)^{-1}<k^{-1}$. On the other hand, as $\kappa$ goes from zero to negative, the power spectrum gets a blue tilt, scaling as $(k/(aH_{\rm inf}))^{-2\kappa/3}$, and then all superhorizon modes are suppressed by the end of inflation.
@@
where the angle bracket denotes the average over different modes and the upper limit of the integral is some ${\cal O}(1)$ number. As we can see from Eq. (II.2), for a positive effective curvature ($\nu>3/2$), the spectrum is indeed red. In this case, the energy density is dominated by the minimum mode $k_{\rm min}$ and the upper limit of the integral does not affect the result. The minimum mode is set by the comoving horizon at the beginning of inflation:
@@
where $a_{\rm i}$ and $\tau_{\rm i}$ are the scale factor and conformal time at the beginning of inflation, respectively. Modes with comoving momentum smaller than $k_{\rm min}$ do not enter the horizon at any time and do not contribute to physical observables. Working out the integral, we obtain the energy density at the end of inflation:
@@ III Relic abundance
Figure 2: Constraints on the axion couplings to the photon (left) and the gluon (right). The couplings are given in Eq. (22). The QCD axion falls along the black line; we dot the line for $m\gtrsim 10^{-4}$ eV, the maximum mass our mechanism can achieve for $N=60$ e-folds of inflation (see Fig. 1). The gray region is excluded by bounds on isocurvature perturbations for $N=60$ e-folds, but this bound is relaxed for more e-folds. We also show a projection Feix et al. (2019) for future isocurvature bounds from CMB-S4 André et al. (2014); Abazajian et al. (2016) and SKA2 Pritchard et al. (2015); Weltman et al. (2020) with the dashed gray line. The green line corresponds to the misalignment production of ALP DM, while the blue region is excluded by the naturalness bound in Eq. (23). In the left panel we present astrophysical bounds on photon-axion conversion Reynés et al. (2021); Reynolds et al. (2020); Marsh et al. (2017); Dessert et al. (2020); Hoof and Schulz (2023); Ajello et al. (2016); Dessert et al. (2022); Noordhuis et al. (2023); Ayala et al. (2014) in purple, bound from axion haloscopes Asztalos et al. (2010); Du et al. (2018); Braine et al. (2020); Bartram et al. (2021); De Panfilis et al. (1987); Hagmann et al. (1990); Zhong et al. (2018); Backes et al. (2021); Jewell et al. (2023); Chang et al. (2022); Quiskamp et al. (2022, 2024); Lee et al. (2020); Jeong et al. (2020); Kwon et al. (2021); Lee et al. (2022); Kim et al. (2023a); Yi et al. (2023); Yang et al. (2023); Kim et al. (2023b); Adair et al. (2022) in red, and projections for future haloscopes as a dashed red line, all adapted from O’Hare (2020). In the right panel we show bounds from supernova 1987a Raffelt (2008) in purple and oscillating neutron EDM experiments Abel et al. (2017) in orange, as well as a projection for a ${}^{229}{\rm Th}$ nuclear clock as a dashed red line, adapted from Kim and Perez (2024).
@@ III.1 Axion-like particles
where $T_{\rm reh}$ is the reheating temperature, and for simplicity we assume an instantaneous reheating after inflation. Due to the $e^{N}$ enhancement on the right-hand side of Eq. (18), it is easy for the axion to become nonrelativistic before structure formation, $T\sim{\rm keV}$.
@@
Note that in this case, the relic abundance does not depend on the axion mass, and it corresponds to a scenario with a much lower inflation scale $H_{\rm inf}$.
@@ III.2 QCD axion
Note that for the QCD axion, the breaking scale $f_{\eta}$ is related to the mass through $f_{\eta}m_{\eta}\approx\Lambda_{\rm QCD}^{2}$. Requiring that $f_{\eta}>H_{\rm inf}/2\pi$ (so that PQ symmetry is broken during inflation) then puts an upper bound on the QCD axion mass, which we present as black curves in Fig. 1. Combined with the backreaction constraint this implies a maximum QCD axion mass of order $10^{-4}$ eV, although this can be relaxed by increasing the number of e-folds of inflation. Compared with the minimal misalignment mechanism without fine-tuning, where the QCD axion mass should be of order $10^{-6}~{}{\rm eV}$ to comprise all of the DM, our mechanism allows for a heavier QCD axion, corresponding to a larger coupling to SM particles. Lastly, in Fig. 1 we also show the lower bound on the QCD axion mass from the requirement that $f_{\eta}<M_{\rm Pl}$.
@@ IV Phenomenology
There are a variety of astrophysical and cosmological bounds on the DM mass that exclude DM lighter than about $10^{-22}$$10^{-20}$ eV, including constraints from the Lyman $\alpha$-forest Rogers and Peiris (2021), pulsar timing array Afzal et al. (2023), star cluster observations Marsh and Niemeyer (2019), the Milky Way mass function Nadler et al. (2021), and black hole superradiance Arvanitaki and Dubovsky (2011) (for a review see e.g. Antypas et al. (2022)). For clarity we have not presented these bounds in Fig. 2.
@@
assuming an $\mathcal{O}(1)$ misalignment angle. The green line in Fig. 2 corresponds to where misalignment yields $\Omega_{\eta}h^{2}\sim 0.12$. Unless one fine-tunes the misalignment angle to suppress this contribution to the relic density, our mechanism is only viable above this line. This is different than the usual misalignment story, where the ALP must lie on this line (or below, with fine-tuning).
Our mechanism allows for a lower PQ breaking scale than misalignment, which means the couplings to SM particles can be larger and easier to detect experimentally. In Fig. 2 we present existing bounds on the axion-photon coupling from astrophysical limits on photon-axion conversion Reynés et al. (2021); Reynolds et al. (2020); Marsh et al. (2017); Dessert et al. (2020); Hoof and Schulz (2023); Ajello et al. (2016); Dessert et al. (2022); Noordhuis et al. (2023); Ayala et al. (2014) and from axion haloscopes Asztalos et al. (2010); Du et al. (2018); Braine et al. (2020); Bartram et al. (2021); De Panfilis et al. (1987); Hagmann et al. (1990); Zhong et al. (2018); Backes et al. (2021); Jewell et al. (2023); Chang et al. (2022); Quiskamp et al. (2022, 2024); Lee et al. (2020); Jeong et al. (2020); Kwon et al. (2021); Lee et al. (2022); Kim et al. (2023a); Yi et al. (2023); Yang et al. (2023); Kim et al. (2023b); Adair et al. (2022), adapted from O’Hare (2020). For the axion-gluon coupling we show bounds from supernova 1987a Raffelt (2008) and from neutron EDM oscillation experiments Abel et al. (2017), adapted from Kim and Perez (2024). There is a wide swath of viable parameter space stretching across a large range of masses, as shown by the white regions in Fig. 2. We include projections for future axion haloscopes (adapted from O’Hare (2020)), such as DANCE Michimura et al. (2020), SRF Berlin et al. (2021), DM-Radio Brouwer et al. (2022), twisted anyon cavity Bourhill et al. (2023) etc., which will effectively probe the photon coupling parameter space. We also show a projection for a future ${}^{229}{\rm Th}$ nuclear clock Kim and Perez (2024) in Fig. 2, which can probe the gluon coupling parameter space.

### 15 Mar 2024 v2

@@ Abstract
We propose a new mechanism to produce axion dark matter from inflationary fluctuations. Quantum fluctuations during inflation are strengthened by a coupling of the axion kinetic term to the inflaton, which we parametrize as an effective curvature $\kappa$ in the axion equation of motion. A nonvanishing curvature breaks the scale invariance of the axion power spectrum, driving a quantum phase transition with $\kappa$ as the order parameter. The axion power spectrum is proportional to the inverse comoving horizon to the power of $\kappa$. For positive $\kappa$ the spectrum gets a red tilt, leading to an exponential enhancement of the axion abundance as the comoving horizon shrinks during inflation. This enhancement allows sufficient axion production to comprise the entire dark matter relic abundance despite the ultralight mass. Our mechanism predicts a significantly different parameter space from the usual misalignment mechanism. It allows for axion-like particle dark matter with a much lower decay constant and thus a larger coupling to Standard Model particles. Much of the parameter space can be probed by future experiments including haloscopes, nuclear clocks, CASPEr, and CMB-S4. We can also generate heavier QCD axion dark matter than the misalignment mechanism.
@@ I Introduction
The usual misalignment mechanism for producing the axion relic abundance predicts a relationship between the axion mass and the PQ breaking scale Dine and Fischler (1983); Preskill et al. (1983); Arias et al. (2012). Assuming that the axion comprises all of the DM, this restricts the viable parameter space. For instance, for QCD axion DM, the mass should be of order $10^{-6}~{}{\rm eV}$ (assuming the initial misalignment angle is not fine-tuned) Grilli di Cortona et al. (2016). On the other hand, if we consider ALP DM lighter than $10^{-12}~{}{\rm eV}$, which is the regime where future quantum sensor probes are most sensitive Kim and Perez (2024); Flambaum and Samsonov (2023); Flambaum and Mansour (2023); Zhao et al. (2024), then the PQ breaking scale has to be higher than $10^{14}~{}{\rm GeV}$. Consequently, the couplings to Standard Model (SM) particles are highly suppressed and difficult to probe experimentally. This has motivated the development of new mechanisms for producing axion DM that allow for a different relationship between mass and coupling, such as parametric resonance Co et al. (2018), nonstandard cosmologies Blinov et al. (2019), kinetic misalignment Co et al. (2020), ALP cogenesis Co et al. (2021), and trapped misalignment Di Luzio et al. (2021a) (see Takahashi et al. (2018); Co et al. (2019); Ho et al. (2019); Harigaya and Leedom (2020); Asadi et al. (2022); Kitano et al. (2023) for more possibilities).
@@
In this work, we propose a new mechanism to produce axion DM from inflationary fluctuations, which allows for a much lower breaking scale than misalignment. The basic idea is to make use of quantum fluctuations during the cosmic inflationary epoch to efficiently produce the axion abundance. Typically, inflationary fluctuations do not lead to a sufficient DM production, due to the smallness of the axion mass Graham et al. (2016); Redi and Tesi (2023). However, if the axion kinetic term is coupled to the inflaton, the relic abundance can be exponentially enhanced [^1] A similar idea was first raised in Ref. Nakai et al. (2020) to produce dark photon DM through exponential enhancement, where a specific form of kinetic coupling was used.. The kinetic coupling breaks the scale-invariant power spectrum of a massless axion and, together with the quantum fluctuations during inflation, drives a “quantum phase transition” from the conformal symmetric phase to the broken phase. We will show how one can calculate the tilt in the power spectrum from the form of the kinetic coupling. We find that if the power spectrum is broken to a red tilt, the axion abundance gets exponentially enhanced by the number of inflationary e-folds, compensating the suppression of the DM density by the small axion mass.
@@
Our mechanism is applicable to both the QCD axion and ALPs. The relic density does not depend on the PQ breaking scale, which opens up a large window in the axion DM parameter space across a wide range of masses down to $\sim 10^{-24}~{}{\rm eV}$. The axion couplings can be much larger than one obtains in the misalignment mechanism. Much of the parameter space can be probed by future axion DM experiments such as haloscopes, nuclear clocks, and CASPEr (see Fig. 3). Our mechanism can also potentially be probed by next-generation cosmic microwave background (CMB) experiments and $21$ cm telescopes.
@@ II.1 Framework
where $M_{\rm Pl}$ is the reduced Planck mass, $R$ is the Ricci scalar, and $V(\phi)$ is the inflaton potential. The axion $\eta$ is coupled to the inflaton $\phi$ through its kinetic term during inflation, with $K(\phi)$ an arbitrary function of the inflaton. Such a noncanonical scalar kinetic term can be naturally realized, e.g., in the framework of supergravity Ellis et al. (1984, 2013) with the scalar $\eta$ identified as the modulus field coming from orbifold compactifications Kobayashi et al. (2016); Abe et al. (2023). An example of PQ unification is given in Appendix A, where a nonminimal kinetic coupling naturally arises. The kinetic function $K(\phi)$ should smoothly reduce to unity at the end of inflation, such that the axion is decoupled from the inflaton thereafter and we are left with a canonical kinetic term. We take the usual flat FLRW metric $g_{\rm\mu\nu}=a^{2}(\tau){\rm diag}\left(-1,1,1,1\right)$, where $a$ is the scale factor and $\tau$ is the conformal time.
@@ II.2 Breaking conformal symmetry
Here $H_{\rm inf}/2\pi$ is the Gibbons–Hawking temperature Gibbons and Hawking (1977) that characterizes the typical magnitude of quantum fluctuations during inflation. In the last step of Eq. (7), we have assumed $\left|\kappa\right|<1$ due to the backreaction constraint (see below).
@@
In particular, a positive $\kappa$ leads to a red tilt, with the axion power spectrum dominated by superhorizon modes. Note that in Eq. (7), we have $1/x=aH_{\rm inf}/k$, which means the power spectrum obeys a power law proportional to the inverse comoving horizon $aH_{\rm inf}$ to the power of $\kappa$. Since the comoving horizon is exponentially shrinking during inflation, the power spectrum will get exponentially enhanced if $\kappa>0$. For each mode $k$, it starts to grow only after exiting the horizon, i.e., $\left(aH_{\rm inf}\right)^{-1}<k^{-1}$. This corresponds to the shaded region in Fig. 1. Modes with smaller $k$ get more enhancement during inflation (larger shaded area). On the other hand, as $\kappa$ goes from zero to negative, the power spectrum gets a blue tilt, scaling as $(k/(aH_{\rm inf}))^{-2\kappa/3}$, and then all superhorizon modes are suppressed by the end of inflation.
Figure 1: A sketch of our mechanism. The comoving horizon shrinks during inflation and increases later. For each mode $k$, with a positive effective curvature, it starts to grow after exiting the horizon (red shaded region). The minimum mode $k_{\rm min}\propto a_{\rm i}H_{\rm inf}$ (red dashed line) is set by the initial horizon, and receives largest enhancement during inflation. The comoving Compton wavelength $1/(am_{\eta})$ is shown with blue line for a fixed axion mass $m_{\eta}$. The mode $k$ becomes nonrelativistic after $a_{\rm NR}(k)=k/m_{\eta}$ and begins to oscillate when $a>a_{\rm osc}$ with $a_{\rm osc}$ determined by $H(a_{\rm osc})=m_{\eta}$. Unlike misalignment, our mechanism allows axion to become nonrelativistic well before coherent oscillation, i.e., $a_{\rm NR}\ll a_{\rm osc}$.
@@
where the angle bracket denotes the average over different modes and the upper limit of the integral is some ${\cal O}(1)$ number. As we can see from Eq. (II.2), for a positive effective curvature ($\nu>3/2$), the spectrum is indeed red. In this case, the upper limit of the integral in (II.2) does not affect the result; the energy density is dominated by the minimum mode $k_{\rm min}$ (see the red dashed line in Fig. 1):
@@
where $a_{\rm i}$ and $\tau_{\rm i}$ are the scale factor and conformal time at the beginning of inflation, respectively. This mode exits the horizon at $\tau_{\rm i}$ and receives the largest exponential enhancement during inflation. Modes with comoving momentum smaller than $k_{\rm min}$ do not enter the horizon at any time and do not contribute to physical observables. Working out the integral, we obtain the energy density at the end of inflation:
@@ III.1 Axion-like particles
where $T_{\rm reh}$ is the reheating temperature, and for simplicity we assume an instantaneous reheating after inflation. Due to the $e^{N}$ enhancement on the right-hand side of Eq. (18), it is easy for the axion to become nonrelativistic before structure formation, $T\sim{\rm keV}$. Therefore, unlike misalignment, our mechanism allows axion to become nonrelativistic much earlier than it begins to oscillate, as can be seen clearly from Fig. 1.
@@
Note that in this case, the relic abundance does not depend on the axion mass, and it corresponds to a scenario with a much lower inflationary Hubble scale $H_{\rm inf}$.
@@
Figure 3: Constraints on the axion couplings to the photon (left) and the gluon (right). The couplings are given in Eq. (22). The QCD axion falls along the black line; we dot the line for $m\gtrsim 10^{-4}$ eV, the maximum mass our mechanism can achieve for $N=60$ e-folds of inflation (see Fig. 2). The gray region is excluded by bounds on isocurvature perturbations for $N=60$ e-folds, but this bound is relaxed for more e-folds. We also show a projection Feix et al. (2019) for future isocurvature bounds from CMB-S4 André et al. (2014); Abazajian et al. (2016) and SKA2 Pritchard et al. (2015); Weltman et al. (2020) with the dashed gray line. The green line corresponds to the misalignment production of ALP DM, while the blue region is excluded by the naturalness bound in Eq. (23). In the left panel we present astrophysical bounds on photon-axion conversion Reynés et al. (2021); Reynolds et al. (2020); Marsh et al. (2017); Dessert et al. (2020); Hoof and Schulz (2023); Ajello et al. (2016); Dessert et al. (2022); Noordhuis et al. (2023); Ayala et al. (2014); Balkin et al. (2020) in purple, bound from axion haloscopes Asztalos et al. (2010); Du et al. (2018); Braine et al. (2020); Bartram et al. (2021); De Panfilis et al. (1987); Hagmann et al. (1990); Zhong et al. (2018); Backes et al. (2021); Jewell et al. (2023); Chang et al. (2022); Quiskamp et al. (2022, 2024); Lee et al. (2020); Jeong et al. (2020); Kwon et al. (2021); Lee et al. (2022); Kim et al. (2023a); Yi et al. (2023); Yang et al. (2023); Kim et al. (2023b); Adair et al. (2022) in red, and projections for future haloscopes O’Hare (2020) as a dashed red line. In the right panel we show astrophysical bounds from supernova 1987a Raffelt (2008); Caputo and Raffelt (2024) and white dwarf cooling Hook and Huang (2018); Balkin et al. (2022) in purple, as well as bounds from oscillating neutron EDM experiments Abel et al. (2017); Schulthess et al. (2022) in orange. We present projections for a ${}^{229}{\rm Th}$ nuclear clock Kim and Perez (2024) and CASPEr-electric Budker et al. (2014); Jackson Kimball et al. (2020) as dashed red and brown lines, respectively. All limits are adapted from O’Hare (2020); Kim and Perez (2024).
@@ III.2 QCD axion
Note that for the QCD axion, the breaking scale $f_{\eta}$ is related to the mass through $f_{\eta}m_{\eta}\approx\Lambda_{\rm QCD}^{2}$. Requiring that $f_{\eta}>H_{\rm inf}/2\pi$ (so that PQ symmetry is broken during inflation) then puts an upper bound on the QCD axion mass, which we present as black curves in Fig. 2. Combined with the backreaction constraint this implies a maximum QCD axion mass of order $10^{-4}$ eV for $N=60$ e-folds, although this can be relaxed by increasing the number of e-folds of inflation. Compared with the minimal misalignment mechanism without fine-tuning, where the QCD axion mass should be of order $10^{-6}~{}{\rm eV}$ to comprise all of the DM, our mechanism allows for a heavier QCD axion, corresponding to a larger coupling to SM particles. Lastly, in Fig. 2 we also show the lower bound on the QCD axion mass from the requirement that $f_{\eta}<M_{\rm Pl}$.
@@ IV Phenomenology
There are a variety of astrophysical and cosmological bounds on the DM mass that exclude DM lighter than about $10^{-22}$$10^{-20}$ eV, including constraints from the Lyman $\alpha$-forest Rogers and Peiris (2021), pulsar timing array Afzal et al. (2023), star cluster observations Marsh and Niemeyer (2019), the Milky Way mass function Nadler et al. (2021), and black hole superradiance Arvanitaki and Dubovsky (2011); Mehta et al. (2020) (for a review see e.g. Antypas et al. (2022)). For clarity we have not presented these bounds in Fig. 3.
@@
assuming an $\mathcal{O}(1)$ misalignment angle, where $h$ is the reduced Hubble constant. The green line in Fig. 3 corresponds to where misalignment yields $\Omega_{\eta}h^{2}\sim 0.12$. Unless one fine-tunes the misalignment angle to suppress this contribution to the relic density, our mechanism is only viable above this line. This is different than the usual misalignment story, where the ALP must lie on this line (or below, with fine-tuning).
Our mechanism allows for a lower PQ breaking scale than misalignment, which means the couplings to SM particles can be larger and easier to detect experimentally. In Fig. 3 we present existing bounds on the axion-photon coupling from astrophysical limits on photon-axion conversion Reynés et al. (2021); Reynolds et al. (2020); Marsh et al. (2017); Dessert et al. (2020); Hoof and Schulz (2023); Ajello et al. (2016); Dessert et al. (2022); Noordhuis et al. (2023); Ayala et al. (2014); Balkin et al. (2020) and from axion haloscopes Asztalos et al. (2010); Du et al. (2018); Braine et al. (2020); Bartram et al. (2021); De Panfilis et al. (1987); Hagmann et al. (1990); Zhong et al. (2018); Backes et al. (2021); Jewell et al. (2023); Chang et al. (2022); Quiskamp et al. (2022, 2024); Lee et al. (2020); Jeong et al. (2020); Kwon et al. (2021); Lee et al. (2022); Kim et al. (2023a); Yi et al. (2023); Yang et al. (2023); Kim et al. (2023b); Adair et al. (2022), adapted from O’Hare (2020). There is a wide swath of viable parameter space stretching across a large range of masses, as shown by the white regions in Fig. 3. We include projections for future axion haloscopes (adapted from O’Hare (2020)), such as DANCE Michimura et al. (2020), SRF Berlin et al. (2021), DM-Radio Brouwer et al. (2022), twisted anyon cavity Bourhill et al. (2023) etc., which will effectively probe the photon coupling parameter space. For the axion-gluon coupling we show astrophysical bounds Raffelt (2008); Caputo and Raffelt (2024); Hook and Huang (2018); Balkin et al. (2022) and constraints from neutron EDM oscillation experiments Abel et al. (2017); Schulthess et al. (2022), adapted from O’Hare (2020); Kim and Perez (2024). The gluon coupling is heavily constrained for masses smaller than $10^{-11}$ eV. Of course, our framework does not require this coupling to be present for an ultralight ALP, as the coefficient of the $\eta G\tilde{G}$ operator is a free parameter of the theory (the generic coupling induced by gravity is Planck-suppressed). The existing constraints simply mean that this type of coupling is disfavored; one could instead consider an ALP that only couples to the photon at tree-level, for instance. We also show projections for a future ${}^{229}{\rm Th}$ nuclear clock Kim and Perez (2024) and for CASPEr-electric Budker et al. (2014); Jackson Kimball et al. (2020) in Fig. 3, which can further probe/constrain the gluon coupling parameter space.